# The Great Ban: Efficacy and Unintended Consequences of a Massive Deplatforming Operation on Reddit

### v1

@@ Conclusions
The Great Ban was a massive deplatforming operation enforced to shut toxic communities on Reddit. To evaluate its effectiveness, we analyzed 16M comments shared over the course of 14 months by 17K users affected by the ban. Our results reveal that 15.6% of the affected users abandoned Reddit after the ban and that those who remained reduced their toxicity by 6.6% on average. Despite this modest toxicity reduction, 5% of users increased their toxicity by more than 70% of their pre-ban level. The presence of such resentful users was widespread across the analyzed subreddits. However, these users likely had little impact on the platform given that only 16% of users increased both toxicity and activity. Likewise, only 21% users increased toxicity and also obtained positive engagement from the community.
@@
Acknowledgments
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity) and by the Italian Ministry of Education and Research (MIUR) in the framework of the CrossLab and the FoReLab projects (Departments of Excellence).

### v5

@@ 6. Conclusions
The Great Ban was a massive deplatforming operation enforced to shut toxic communities on Reddit. To evaluate its effectiveness, we analyzed 16M comments shared over the course of 14 months by 17K users affected by the ban. Our results reveal that 15.6% of the affected users abandoned Reddit after the ban and that those who remained reduced their toxicity by 6.6% on average. Despite this modest toxicity reduction, 5% of users increased their toxicity by more than 70% of their pre-ban level. The presence of such resentful users was widespread across the analyzed subreddits rather than concentrated in a few ones.
@@
Acknowledgements.
This work is partially supported by the project PIANO (Personalized Interventions Against Online Toxicity) under CUP B53D23013290006 as part of the Italian PRIN 2022, and funded by the European Union’s program Next Generation EU; by the PNRR-M4C2 (PE00000013) "FAIR-Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded under Next Generation EU; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).